OpenStudy (anonymous):

Find the constant of variation for the quadratic variation. 6y = 9x2

OpenStudy (cggurumanjunath):

y=kx^2

OpenStudy (cggurumanjunath):

y=9/6* x^2

OpenStudy (cggurumanjunath):

k=constant of variation

OpenStudy (cggurumanjunath):

k=9/6=3/2=1.5

OpenStudy (anonymous):

I have no clue what that means

OpenStudy (cggurumanjunath):

the quadratic variation is of the form below. y=kx^2

OpenStudy (anonymous):

Thank you

OpenStudy (cggurumanjunath):

:)

OpenStudy (anonymous):

is it 3/2?

OpenStudy (cggurumanjunath):

yes.

OpenStudy (anonymous):

Thank you

OpenStudy (cggurumanjunath):

Anytime.